# Topology-Enhanced Superconducting Qubit Networks for In-Sensor Quantum Information Processing

### v1

@@ 1 Introduction
The paper is thus organized as follows. After the introduction of the fundamental physics of the isolated flux qubit (Sec. II) and of flux qubit networks (Sec. III), we investigate in Sec. IV the static properties of arrays in two distinct geometries (i.e., LA and CA), emphasizing the critical role played by topology in shaping the pattern of circulating currents. The magnetic flux generated by arrays with different topological configurations and its potential detection are issues addressed in Sec. V. In Sec. VI, we turn to the analysis of qubit arrays coupled to a transmission line. By combining exact diagonalization techniques with linear response theory, we provide a detailed characterization of the dynamical response of the arrays to external perturbations. The analysis of the arrays response unambiguously demonstrates that both spectral and dynamical properties are strongly governed by the underlying topological structure, offering key insights for the design and optimization of topologically-engineered quantum devices. Finally, in Sec. VII, we test the expressive capabilities of a quantum reservoir computer based on a superconducting qubit network in a prototypical prediction task and present our conclusions in Sec. VIII.
@@ 2 Physics of an Isolated Flux Qubit
As a preliminary step towards the analysis of qubit arrays, we review the fundamental properties of an isolated flux qubit (Fig. 1(a)), introducing the relevant notation and key flux-dependent quantities that will be used throughout. The quantum dynamics of a superconducting flux qubit can be described by an effective Hamiltonian of the form[7]:
@@
[…]
@@
[…]
@@
(|±$(|\uparrow\rangle\pm|\downarrow\rangle)/\sqrt{2},$ (11)
@@
The externally applied magnetic flux $\Phi_{\mathrm{ex}}$ acts as a tunable control parameter for the flux qubit, governing the direction of the persistent supercurrent circulating in the superconducting loop. This behavior becomes evident when analyzing the expectation value of the ground-state current as a function of $\Phi_{\mathrm{ex}}$ (Fig. 1(d)). In response to an applied flux, the system establishes a circulating current that adjusts the total magnetic flux threading the loop toward quantized values, either 00 or $\Phi_{0}$, corresponding to the minima of an effective double-well potential.
@@ 4 Spectral Properties and Loop-Current Distribution of Linear and Cross-Shaped Arrays
We analyze the energy spectrum and loop-current distribution of two distinct five-qubit superconducting networks, namely a linear array and a cross-shaped arrangement, whose qubits are coupled exclusively via mutual inductance. Because a counterclockwise loop current in one qubit generates a negative magnetic flux in its neighbors, the mutual inductance coefficients are taken as negative ($M_{ij}<0$). Couplings beyond nearest neighbors are neglected, since the magnetic interaction between distant loops only provides small corrections. Throughout the simulations, the reduced magnetic flux is fixed at $f=0.52$, i.e., slightly above the optimal working point $f=1/2$ of the qubit.
Figure 3(a) shows the energy spectrum of the linear array obtained by exact diagonalization. The spectrum exhibits a non-degenerate energy-level structure (consistent with the one-dimensional irreducible representations of the $C_{2}$ symmetry group), with the lowest states well separated from the excited states, reflecting the collective nature of the excitations. In the ground state, the loop currents $\langle\hat{I}^{(i)}\rangle$ are mirror symmetric with respect to the center of the chain [Fig. 3(b)], with the maximal current amplitude observed in the central qubit. This distribution is a direct consequence of the array topology and the nature of the inductive couplings. The current profile associated with the first excited state displays a similar pattern, even though the loop currents exhibit clockwise circulation and reduced magnitude.
@@
The energy spectrum of the cross-shaped array is reported in Fig. 4(a). In this configuration, qubit 2 occupies the central position and is inductively coupled to the four peripheral qubits. Due to the $C_{4v}$ point-group symmetry (90 rotations and mirror lines) of the array, the low-energy sector of the energy spectrum organizes into two groups of nearly generate levels (corresponding to states presenting both accidental quasi-degeneracy and symmetry-protected degeneracy), which play a relevant role in defining an evident energy gap between the first excited state and the second excited state. This symmetry-induced phenomenon, not observed in the linear array case, appears to be relevant in shaping the response of a quantum device with function-oriented design. Moreover, the enhanced connectivity of this geometry significantly alters the current distribution. Indeed, as shown in Fig. 4(b), the ground-state current profile is now highly localized at the central qubit, which carries the largest loop current. This behavior is also observed in the first excited state, where the current enhancement factor is even more pronounced compared to the one observed in the ground state.
@@
Figure 3: (a) Energy eigenvalues of a linear array of five identical qubits. Energy eigenvalues are expressed in units of $\hbar\omega_{0}=I_{s}\Phi_{0}$. (b) Loop currents of the linear array normalized to the maximum current $I_{s}$ as a function of the qubit index $i$ for the ground state and the first excited state. (c) Current-current correlations, in units of $I_{S}^{2}$, between the first qubit and the $i$-th qubit.
@@
The marked difference between the low-energy spectra of the linear and cross-shaped arrays highlights the crucial role that network topology plays in determining the emergent properties of the system. In particular, the cross-shaped array exhibits a sort of quantum amplification of the central loop current, similar to the localization phenomenon described in Ref. [40]. Interestingly, the latter can be interpreted as a cooperative effect induced by network connectivity and can also be understood in terms of concepts of spectral graph theory. In graph theory terms, the nearest-neighbor inductive couplings define the adjacency matrix of an undirected graph in which each qubit represents a node. Hubs in this network, i.e. qubits sharing a high number of inductive interactions, play a fundamental role in defining the quantum-network properties.
@@
Another promising perspective consists of encoding a logical qubit in the many-body manifold spanned by the collective ground state $\lvert G\rangle$ and the first excited state $\lvert E\rangle$ of the entire cross-shaped array. In view of the presented spectral properties, this choice may passively suppress uncorrelated decoherence through symmetry-protected subradiance. The expected advantage over a single qubit complemented by active error correction must be validated experimentally by charting the regimes of temperature, fabrication tolerance, and noise correlation where this protection proves effective.
@@
Figure 4: (a) Energy eigenvalues of a cross-shaped array of five identical qubits. Energy eigenvalues are expressed in units of $\hbar\omega_{0}=I_{s}\Phi_{0}$. (b) Loop currents of the cross-shaped array normalized to the maximum current $I_{s}$ as a function of the qubit index $i$ for the ground state and the first excited state.
@@ 6 Response of a Qubit Network to Dynamical Coupling with a Transmission Line
where $H_{0}$ represents the Hamiltonian of the superconducting qubit network. The coupling term with the transmission line can be expressed as $H_{I}(t)=f(t)\ \theta(t-t_{0})\ B_{S}$, where $f(t)$ is a time-dependent function proportional to the current $I(t)$ flowing in the transmission line, $B_{S}$ is an appropriate operator in Schrödinger representation, while the Heaviside step function $\theta(t-t_{0})$ ensures that the coupling is activated only after the onset time $t_{0}$. The expectation value of a generic operator $A_{S}$ in Schrödinger representation, purified from contributions not related to the external perturbation, is given by the convolution integral:
[…]
@@ 6.1 Dynamical Response of a Linear Array
Figure 6: (a) Response function amplitude $|\chi(\omega)|$ of a linear array of five inductively coupled qubits, plotted versus the normalized drive frequency $\omega/\omega_{0}$ (model parameters: $\Phi_{ex}/\Phi_{0}=0.52$, $\Delta=-MI_{S}^{2}=0.2\ \hbar\omega_{0}$, $\eta=2.5\cdot 10^{-3}\ \hbar\omega_{0}$, with $\omega_{0}=I_{s}\Phi_{0}/\hbar$). (b) Corresponding response function phase $\phi(\omega)/\pi$. (c,d) Same quantities for the uncoupled case (mutual inductance coupling energy $-MI_{s}^{2}$ set to $10^{-6}\ \hbar\omega_{0}$), i.e. five isolated qubits. Comparison of (a,b) with (c,d) highlights the emergence of multiple resonances and phase discontinuities induced by the qubit–qubit coupling, in contrast to the single resonance and simple $\pi$–phase roll exhibited by uncoupled qubits.
@@
In contrast, panels (c) and (d) report the amplitude and phase response for five isolated qubits (with $-MI_{s}^{2}=10^{-6}\ \hbar\omega_{0}$). In this uncoupled limit, the amplitude spectrum in (c) reduces to a single Lorentzian resonance at $\omega/\omega_{0}\approx 0.40$, and the phase in (d) displays the textbook $\pi$ phase roll-off characteristic of an isolated two‐level system. The direct comparison of (a,b) with (c,d) highlights the emergence of multiple resonances and phase discontinuities as unambiguous signatures of the qubit-qubit inductive coupling.
@@ 6.3 Dynamical Response of Disordered Arrays
So far, we have considered arrays of identical flux qubits, characterized by uniform parameters. However, in realistic implementations, fabrication imperfections inevitably introduce random variations in qubit properties [37]. To assess the robustness of the system’s response under such imperfections, we investigate the effects of disorder in a linear array of flux qubits, which serves as a relevant study case to be compared to the response of the ideal system presented in Sec. 6.1. Disorder is incorporated in the model by replacing the loop critical currents $I_{S}^{(i)}$ and the tunneling energies $\Delta^{(i)}$ with $I_{S}(1+\lambda_{i})$ and $\Delta(1+\mu_{i})$, respectively, where $\lambda_{i}$ and $\mu_{i}$ are independent random variables uniformly distributed in the interval $[-0.1,0.1]$. This choice reflects fabrication-induced parameter fluctuations up to ±$\pm 10\ \%$ of the nominal values $I_{s}$ and $\Delta$. These modifications affect both the mutual inductive coupling between qubits and their interaction with the transmission line. In particular, the mutual coupling between qubits $i$ and $j$ is now weighted by the extra-factor $(1+\lambda_{i})(1+\lambda_{j})$, while the coupling to the transmission line takes the modified form:
@@
Figure 9: (a) $|\chi(\omega)|$ of a disordered linear array of five coupled qubits, plotted versus the normalized drive frequency $\omega/\omega_{0}$. The system parameters are: $\Phi_{ex}/\Phi_{0}=0.52$, $I_{S}^{(i)}=I_{s}(1+\lambda_{i})$, $\Delta_{i}=\Delta(1+\mu_{i})$, $\Delta=-MI_{S}^{2}=0.2\ \hbar\omega_{0}$, $\eta=2.5\cdot 10^{-3}\ \hbar\omega_{0}$, with $\omega_{0}=I_{s}\Phi_{0}/\hbar$. The random variables $\lambda_{i}$ and $\mu_{i}$ have been obtained as detailed in the main text. (b) Corresponding response function phase $\phi(\omega)/\pi$. (c,d) $|\chi(\omega)|$ and $\phi(\omega)/\pi$ for five isolated qubits, i.e. a system with mutual inductance coupling energy $-MI_{s}^{2}$ set to $10^{-6}\ \hbar\omega_{0}$, while taking the remaining parameters as in panels (a) and (b). (e) Color–density map of $\lvert\chi(\omega)\rvert$ for a disordered linear qubit array coupled to a transmission line, shown as a function of the normalized drive frequency $\omega/\omega_{0}$ and the applied flux $\Phi_{\mathrm{ex}}/\Phi_{0}$. Model parameters are set as done in panel (a) and (b). (f) Color–density map of $\phi(\omega)/\pi$ under the same conditions as in (e). In all density plots, darker shading corresponds to lower values of the plotted quantity.

### v2

@@ 1 Introduction
The paper is thus organized as follows. After the introduction of the fundamental physics of the isolated flux qubit (Sec. 2) and of flux qubit networks (Sec. 3), we introduce the coherence-assisted topological enhancement paradigm in Sec. 4, and we investigate in Sec. 5 the static properties of arrays in two distinct geometries (i.e., LA and CA), emphasizing the critical role played by topology in shaping the pattern of circulating currents. The magnetic flux generated by arrays with different topological configurations and its potential detection are issues addressed in Sec. 6. In Sec. 7, we turn to the analysis of qubit arrays coupled to a transmission line. By combining exact diagonalization techniques with linear response theory, we provide a detailed characterization of the dynamical response of the arrays to external perturbations. The analysis of the arrays response unambiguously demonstrates that both spectral and dynamical properties are strongly governed by the underlying topological structure, offering key insights for the design and optimization of topologically-engineered quantum devices. Finally, in Sec. 8, we test the expressive capabilities of a quantum reservoir computer based on a superconducting qubit network in a prototypical prediction task and present our conclusions in Sec. 9.
@@ 2 Physics of an Isolated Flux Qubit
As a preliminary step towards the analysis of qubit arrays, we review the fundamental properties of an isolated flux qubit (Fig. 1(a)), introducing the relevant notation and key flux-dependent quantities that will be used throughout. The quantum dynamics of a superconducting flux qubit can be described by an effective Hamiltonian of the form[7, 16, 20, 8, 11]:
@@
Introducing the eigenstates $\{|\uparrow\rangle,|\downarrow\rangle\}$ of the Pauli matrix $\sigma_{z}$, defined by the relations:
@@
From a physics point of view, the state $|\uparrow\rangle$ represents a counterclockwise circulating current, whereas the state $|\downarrow\rangle$ represents a clockwise current (Fig. 1(b)).
Figure 1: (a) Schematic of a flux qubit realized by means of the four Josephson junctions technology [48]. The red crosses denote the weak links between superconducting regions (blue lines). (b) Circulating currents in the qubit loop: counterclockwise and clockwise currents correspond to the current states $|\uparrow\rangle$ and $|\downarrow\rangle$, respectively. (c) Energy spectrum of the effective Hamiltonian $H_{\text{eff}}$ as a function of the normalized magnetic flux $f$. (d) Expectation value of the current operator on the energy eigenstates, normalized to the maximum loop current $I_{s}$ and plotted as a function of $f$.
@@
$(|\uparrow\rangle\pm|\downarrow\rangle)/\sqrt{2},$ (11)
@@
The externally applied magnetic flux $\Phi_{\mathrm{ex}}$ acts as a tunable control parameter for the flux qubit, governing the direction of the persistent supercurrent circulating in the superconducting loop. This behavior becomes evident when analyzing the expectation value of the ground-state current as a function of $\Phi_{\mathrm{ex}}$ (Fig. 1(d)). In response to an applied flux, the system establishes a circulating current that adjusts the total magnetic flux threading the loop toward quantized values, either 0 or $\Phi_{0}$, corresponding to the minima of an effective double-well potential.
@@ 3 Superconducting Qubit Networks
4 Coherence-Assisted Topological Enhancement Paradigm
The present analysis is motivated by a line of research that has progressively unveiled the fundamental role of the interplay of topology and coherence in cold bosonic atoms and in superconducting systems defined on graph-like structures.
Indeed, from a theoretical standpoint, Refs. [38, 39, 40, 41, 42] have shown that inhomogeneous connectivity can induce nontrivial collective phenomena such as Bose–Einstein condensation on star and comb graphs or the emergence of topological focusing in discrete Ginzburg–Landau models. In these systems, nodes with higher degree act as attractive centers for the condensate or for the superconducting order parameter.
These studies demonstrated that the network topology itself can promote spatial localization and amplification phenomena, effectively turning connectivity into a control parameter of the macroscopic quantum state.
Theoretical predictions have subsequently found experimental confirmation in Josephson-junction networks [43, 44, 45, 46], where coherent and topological effects have been directly observed through critical current enhancement and long-range phase correlations. These experiments have established a clear connection between graph topology and superconducting coherence in realistic architectures of connected superconducting islands.
Analogous effects also arise in the discrete Schrödinger equation on graphs, whereas they are absent in purely classical diffusive processes, where the stationary state is homogeneous and the topology affects only the relaxation dynamics, not the final equilibrium configuration. This distinction emphasizes that topological effects are relevant exclusively to wave-like, coherent systems, in which nontrivial interference patterns sustain organized spatial structures.
Building on these theoretical and experimental advances, our goal is to extend the concepts of coherence-assisted topological enhancement to superconducting qubit networks. In this context, we aim to investigate how similar mechanisms can be harnessed within architectures of coupled flux qubits, with potential technological implications for quantum sensing and in-sensor quantum information processing.
To explore the interplay between network topology and coherence, we focus below on two paradigmatic five-qubit architectures: a linear array and a cross-shaped array. From a graph-theoretical perspective, a linear array is described by a tridiagonal adjacency matrix in which the node connectivity varies between one and two, depending on whether the node is terminal or internal. In contrast, a cross-shaped array includes a central node with connectivity four and four peripheral nodes with connectivity one. These distinct topological features are expected to yield markedly different collective behaviors. The purpose of the following analysis is to verify that topological enhancement, previously demonstrated and experimentally confirmed in other superconducting platforms, also manifests in flux-qubit networks.
@@ 5 Spectral Properties and Loop-Current Distribution of Linear and Cross-Shaped Arrays
In line with the program outlined above, we proceed by examining the energy spectrum and loop-current distribution of two distinct five-qubit superconducting networks, namely a linear array and a cross-shaped arrangement, whose qubits are coupled exclusively via mutual inductance. Because a counterclockwise loop current in one qubit generates a negative magnetic flux in its neighbors, the mutual inductance coefficients are taken as negative ($M_{ij}<0$). Couplings beyond nearest neighbors are neglected, since the magnetic interaction between distant loops only provides small corrections. Throughout the simulations, the reduced magnetic flux is fixed at $f=0.52$, i.e., slightly above the optimal working point $f=1/2$ of the qubit.
Figure 3(a) shows the energy spectrum of the linear array obtained by exact diagonalization of the Hamiltonian:
$H_{LA}=-\sum_{i=1}^{5}[\epsilon(f)\sigma_{z}^{(i)}+\Delta\ \sigma_{x}^{(i)}]+MI_{S}^{2}\sum_{i=1}^{4}\sigma_{z}^{(i)}\sigma_{z}^{(i+1)},$ (16)
with model parameters given in the figure inset. The spectrum exhibits a non-degenerate energy-level structure (consistent with the one-dimensional irreducible representations of the $C_{2}$ symmetry group), with the lowest states well separated from the excited states, reflecting the collective nature of the excitations. In the ground state, the loop currents $\langle\hat{I}^{(i)}\rangle$ are mirror symmetric with respect to the center of the chain [Fig. 3(b)], with the maximal current amplitude observed in the central qubit. This distribution is a direct consequence of the array topology and the nature of the inductive couplings. The current profile associated with the first excited state displays a similar pattern, even though the loop currents exhibit clockwise circulation and reduced magnitude.
@@
The energy spectrum of the cross-shaped array, derived from the Hamiltonian:
$H_{CA}=-\sum_{i=1}^{5}[\epsilon(f)\sigma_{z}^{(i)}+\Delta\ \sigma_{x}^{(i)}]+MI_{S}^{2}\ \sigma_{z}^{(2)}\sum_{i\neq 2}\sigma_{z}^{(i)},$ (17)
@@
is reported in Fig. 4(a). In this configuration, qubit 2 occupies the central position and is inductively coupled to the four peripheral qubits. Due to the $C_{4v}$ point-group symmetry (90 rotations and mirror lines) of the array, the low-energy sector of the energy spectrum organizes into two groups of nearly generate levels (corresponding to states presenting both accidental quasi-degeneracy and symmetry-protected degeneracy), which play a relevant role in defining an evident energy gap between the first excited state and the second excited state. This symmetry-induced phenomenon, not observed in the linear array case, appears to be relevant in shaping the response of a quantum device with function-oriented design. Moreover, the enhanced connectivity of this geometry significantly alters the current distribution. Indeed, as shown in Fig. 4(b), the ground-state current profile is now highly localized at the central qubit, which carries the largest loop current. This behavior is also observed in the first excited state, where the current enhancement factor is even more pronounced compared to the one observed in the ground state.
@@
Figure 3: (a) Energy eigenvalues $E_{n}$ of a linear array of five identical qubits, labeled by the index $n$, and sorted in ascending order. Energy eigenvalues are expressed in units of $\hbar\omega_{0}=I_{s}\Phi_{0}$. (b) Loop currents of the linear array normalized to the maximum current $I_{s}$ as a function of the qubit index $i$ for the ground state and the first excited state. (c) Current-current correlations, in units of $I_{S}^{2}$, between the first qubit and the $i$-th qubit.
@@
The marked difference between the low-energy spectra of the linear and cross-shaped arrays highlights the crucial role that network topology plays in determining the emergent properties of the system. In particular, the cross-shaped array exhibits a marked enhancement of the central loop current, consistent with the topological focusing mechanism discussed in Sec. 4 and reminiscent of the localization effects previously discussed within the Ginzburg–Landau framework in Ref. [40]. This amplification can be viewed as a cooperative effect induced by network connectivity and can also be interpreted within the framework of spectral graph theory. In this context, the nearest-neighbor inductive couplings define the adjacency matrix of an undirected graph, in which each qubit represents a node. Highly connected nodes act as hubs that govern the collective response of the system, serving as focal points for magnetic flux concentration and thus shaping the overall properties of the quantum network.
@@
The spectral properties of the cross-shaped array suggest another promising perspective consisting of encoding a logical qubit in the many-body manifold spanned by the collective ground state $\lvert G\rangle$ and the first excited state $\lvert E\rangle$ of the entire cross-shaped array. In view of the presented spectral properties, this choice may passively suppress uncorrelated decoherence through symmetry-protected subradiance. The expected advantage over a single qubit complemented by active error correction must be validated experimentally by charting the regimes of temperature, fabrication tolerance, and noise correlation where this protection proves effective.
@@
Figure 4: (a) Energy eigenvalues $E_{n}$ of a cross-shaped array of five identical qubits, labeled by the index $n$, and sorted in ascending order. Energy eigenvalues are expressed in units of $\hbar\omega_{0}=I_{s}\Phi_{0}$. (b) Loop currents of the cross-shaped array normalized to the maximum current $I_{s}$ as a function of the qubit index $i$ for the ground state and the first excited state.
@@ 7 Response of a Qubit Network to Dynamical Coupling with a Transmission Line
where $H_{0}$, identical to $H$ in Eq. (12), represents the Hamiltonian of the generic flux-qubit network. The coupling term with the transmission line can be expressed as $H_{I}(t)=f(t)\ \theta(t-t_{0})\ B_{S}$, where $f(t)$ is a time-dependent function proportional to the current $I(t)$ flowing in the transmission line, $B_{S}$ is an appropriate operator in Schrödinger representation, while the Heaviside step function $\theta(t-t_{0})$ ensures that the coupling is activated only after the onset time $t_{0}$. The expectation value of a generic operator $A_{S}$ in Schrödinger representation, purified from contributions not related to the external perturbation, is given by the convolution integral:
$\delta\langle A(t)\rangle=\int_{t_{0}}^{\infty}\chi(t-t^{\prime})f(t^{\prime})dt^{\prime},$ (20)
@@ 7.1 Dynamical Response of a Linear Array
Figure 6: (a) Response function amplitude $|\chi(\omega)|$ of a linear array of five inductively coupled qubits, plotted versus the normalized drive frequency $\omega/\omega_{0}$ (model parameters: $\Phi_{ex}/\Phi_{0}=0.52$, $\Delta=-MI_{S}^{2}=0.2\ \hbar\omega_{0}$, $\eta=2.5\cdot 10^{-3}\ \hbar\omega_{0}$, with $\omega_{0}=I_{s}\Phi_{0}/\hbar$). (b) Corresponding response function phase $\phi(\omega)/\pi$. (c,d) Same quantities for the uncoupled case (mutual inductance coupling energy $MI_{s}^{2}$ set to $10^{-6}\ \hbar\omega_{0}$), i.e. five isolated qubits. Comparison of (a,b) with (c,d) highlights the emergence of multiple resonances and phase discontinuities induced by the qubit–qubit coupling, in contrast to the single resonance and simple $\pi$–phase roll exhibited by uncoupled qubits.
@@
In contrast, panels (c) and (d) report the amplitude and phase response for five isolated qubits (with $MI_{s}^{2}=10^{-6}\ \hbar\omega_{0}$). In this uncoupled limit, the amplitude spectrum in (c) reduces to a single Lorentzian resonance at $\omega/\omega_{0}\approx 0.40$, and the phase in (d) displays the textbook $\pi$ phase roll-off characteristic of an isolated two‐level system. The direct comparison of (a,b) with (c,d) highlights the emergence of multiple resonances and phase discontinuities as unambiguous signatures of the qubit-qubit inductive coupling.
@@ 7.3 Dynamical Response of Disordered Arrays
So far, we have considered arrays of identical flux qubits, characterized by uniform parameters. However, in realistic implementations, fabrication imperfections inevitably introduce random variations in qubit properties [37]. To assess the robustness of the system’s response under such imperfections, we investigate the effects of disorder in a linear array of flux qubits, which serves as a relevant study case to be compared to the response of the ideal system presented in Sec. 7.1. Disorder is incorporated in the model by replacing the loop critical currents $I_{S}^{(i)}$ and the tunneling energies $\Delta^{(i)}$ with $I_{S}(1+\lambda_{i})$ and $\Delta(1+\mu_{i})$, respectively, where $\lambda_{i}$ and $\mu_{i}$ are independent random variables uniformly distributed in the interval [0.1,0.1][-0.1,0.1]. This choice reflects fabrication-induced parameter fluctuations up to $\pm 10\ \%$ of the nominal values $I_{s}$ and $\Delta$. These modifications affect both the mutual inductive coupling between qubits and their interaction with the transmission line. In particular, the mutual coupling between qubits $i$ and $j$ is now weighted by the extra-factor $(1+\lambda_{i})(1+\lambda_{j})$, so that the Hamiltonian of the isolated array can be written as:
$H_{LA}^{(D)}=-\sum_{i=1}^{5}[(1+\lambda_{i})\epsilon(f)\sigma_{z}^{(i)}+(1+\mu_{i})\Delta\sigma_{x}^{(i)}]+$
$+MI_{S}^{2}\sum_{i=1}^{4}(1+\lambda_{i})(1+\lambda_{i+1})\sigma_{z}^{(i)}\sigma_{z}^{(i+1)},$ (28)
while the coupling to the transmission line takes the modified form:
@@
Figure 9: (a) $|\chi(\omega)|$ of a disordered linear array of five coupled qubits, plotted versus the normalized drive frequency $\omega/\omega_{0}$. The system parameters are: $\Phi_{ex}/\Phi_{0}=0.52$, $I_{S}^{(i)}=I_{s}(1+\lambda_{i})$, $\Delta_{i}=\Delta(1+\mu_{i})$, $\Delta=-MI_{S}^{2}=0.2\ \hbar\omega_{0}$, $\eta=2.5\cdot 10^{-3}\ \hbar\omega_{0}$, with $\omega_{0}=I_{s}\Phi_{0}/\hbar$. The random variables $\lambda_{i}$ and $\mu_{i}$ have been obtained as detailed in the main text. (b) Corresponding response function phase $\phi(\omega)/\pi$. (c,d) $|\chi(\omega)|$ and $\phi(\omega)/\pi$ for five isolated qubits, i.e. a system with mutual inductance coupling energy $MI_{s}^{2}$ set to $10^{-6}\ \hbar\omega_{0}$, while taking the remaining parameters as in panels (a) and (b). (e) Color–density map of $\lvert\chi(\omega)\rvert$ for a disordered linear qubit array coupled to a transmission line, shown as a function of the normalized drive frequency $\omega/\omega_{0}$ and the applied flux $\Phi_{\mathrm{ex}}/\Phi_{0}$. Model parameters are set as done in panel (a) and (b). (f) Color–density map of $\phi(\omega)/\pi$ under the same conditions as in (e). In all density plots, darker shading corresponds to lower values of the plotted quantity.